OpenStudy (anonymous):

If a cylinder has the dimensions 4 cm as the height, and 2 cm as the radius, what the dimensions of the new cylinder be if the scale factor is 3:1?

OpenStudy (anonymous):

we use a formula

OpenStudy (anonymous):

oh, its just a proportion

OpenStudy (anonymous):

Yes, what is the formula?

OpenStudy (anonymous):

(4x / 2x) = 3/ 1

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

ok thanks!

OpenStudy (anonymous):

wait that dont make sense

OpenStudy (anonymous):

hannnnnnnnahhhhhhh!!!!! you have a new picture

OpenStudy (anonymous):

so pretty

OpenStudy (anonymous):

i not sure about the order here

OpenStudy (anonymous):

ignore what i said before. theres a diferent want to do this

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

New linear measurement = Old linear measurement * scale factor

OpenStudy (anonymous):

ok ?

OpenStudy (anonymous):

So then would the dimensions of the original cylinder be multiplied by the scale factor? I think?

OpenStudy (anonymous):

so we have 4 as an old linear measurement. so to find the new linear measurement NL = 4 * ( 3/1) , where NL is new linear measurement

OpenStudy (anonymous):

Oh ok! Thanks!

OpenStudy (anonymous):

so the answer is 12 , and the other one is 6

OpenStudy (anonymous):

is that the answer in the book

OpenStudy (anonymous):

Yeah, should be thanks